Fermi National Laboratory


Accelerator Update

Wednesday June 9
The day shift began with Operations monitoring the Recycler stash, with Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-zero (D0) in access, and with MiniBooNE taking beam.

Operations turned off the Switchyard beam at 8:32 AM in preparation for shot setup.

Operations put the TeV into shot setup at 8:45 AM. This was the first Mixed-Antiproton shot setup where both the Pbar Accumulator and the Recycler contributed antiprotons to the Tevatron (TeV) store.

At 10:53 AM, a MiniBooNE expert said that beam wasn't making it to their target. Operations called an expert who found a broken switching module. He had it repaired and back online by 11:34 AM.

By 11:46 AM, experts had loaded seven shots of antiprotons from Pbar and two from the Recycler.

Operations established store 3564 at 12:02 PM with an initial luminosity of 66.66E30. A Pbar expert began an antiproton measurement - no stacking.

Pbar resumed stacking at 1:28 PM.

A Linac expert noticed at 5:30 PM that the I- Source had no trigger signal. He power cycled the fiber optic crate and restored normal operations.

At 6:38 PM, Booster suffered an extraction problem. Operators investigated and discovered a kicker (MK04) that needed a timing adjustment. Pbar resumed stacking.

At 6:55 PM, a kicker (MKS08) tripped off. Operators bypassed the kicker and called in an expert to investigate. He found and repaired a failed PFN cable.

Pbar was able to resume stacking at 7:52 PM.

CDF requested at 8:42 PM that Operations fix a colliding beam problem by adjusting collimators. Operators made the adjustments.

At 10:05 PM, CDF asked if Operations could move the Pots away from the beam. Operators contacted a TeV expert and then moved the Pots. CDF turned back on at 10:57 PM.

At 11:27 PM, CDF asked Operations to make more collimator adjustments. Operators called an expert and then made the move.

Thursday June 10
The midnight shift began with Operations monitoring store 3564, with Pbar stacking, and with MiniBooNE taking beam.

Operations made another collimator move at 3:18 AM and another move at 8:07 AM.

Operations turned off MiniBooNE to prepare the area for an access at 9:30 AM.

At 9:56 AM, the Pbar Accumulator injection septum tripped off. An expert investigated, but couldn't find anything wrong. Stacking resumed a few minutes later.

Operations turned off the Main Injector (MI) and Booster at 10 AM.

Linac technicians turned off a RF station (LRF1) at 10:04 AM to repair a LCW leak. This halted all downstream beam.

By 10:30 AM, techs and experts were accessing MI, Booster, Linac, and the Preaccelerator for maintenance and repairs.

All accesses had ended and Operations had beam in MI and to the Accumulator by 4 PM.

A Linac RF station (LRF2) tripped off at 4:45 PM for the first of five times in twenty minutes. It reset okay each time.

A MI RF expert said that MIRF16 would not be available for shots.

LRF2 tripped off again at 6:40 PM. It reset okay.

LRF3 tripped off at 9:20 PM for the first of three times in sixteen minutes. It reset okay each time.

Operations reported at 10:04 PM that the Recycler had suffered a quadrupole magnet failure. An expert came in to investigate.

Operations terminated store 3564 at 10:17 PM.

Operations put the TeV through a dry squeeze at 10:43 PM.

At 10:44 PM, BRF16 tripped off for the first of three times within three minutes. It reset okay each time.

At 11:16 PM, the Pbar sequencer failed due to a "cavity short out" indication. Operators investigated and discovered it had been inadvertently left in that condition during an earlier study.

Operations put the TeV into shot setup at 11:17 PM.

Friday June 11
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations reported at 1:43 AM that they'd been having problems keeping MiniBooNE up and running. Operators had, by 4:10 AM, made some temporary tuning changes that got them back on line.

Operations loaded the protons into the TeV at 3:42 AM, but they aborted fifteen minutes later.

Operations established store 3567 at 6:44 AM with an initial luminosity of 75.17E30.

While a Pbar expert was conducting a study, the remaining 17mA stack was lost at 7:53 AM. Operators investigated, but couldn't find a cause for the stack loss.

Pbar resumed stacking at 8 AM.

Plans for Today
Shortly after shift change the Main Injector developed a ground fault. This needs to be found and fixed. TeV experts will work on fixing the colliding beam problems that are causing interruptions to CDF data taking. This weekend will consist of stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/11/2004   email Fermilab